OpenStudy (anonymous):

Solve by the elimination method 3x-7y=15 -5x+4y=-25

OpenStudy (anonymous):

multiply the top problem by 5 the whole way through and the second one by 3 to elimate the x and then we will solve for y

OpenStudy (anonymous):

15x - 35y = 75 -15x + 12y = -75 ----------------- -23y = 0 y = 0 3x - 0 = 15 x = 5

OpenStudy (anonymous):

therefore you will have: 15x-35y=75 -15x+12y=75 0 -23y=0 y=0 -5x=-25 x=5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

well they text the answer first but i wanted to saw you step by step